OpenStudy (anonymous):

What is the final simplified answer to: square root 32 / square root 2 square root 28 / square root 7

OpenStudy (anonymous):

\[\frac{\sqrt{32}}{\sqrt{2}}=\sqrt{16}=4\]

OpenStudy (anonymous):

\[\frac{\sqrt{28}}{\sqrt{7}}=\sqrt{4}=2\]